consider the formula A=s2 for the area A of a square with a side of length s.
a. graph all possible values of s and a on a coordinate plane.
b. explain how the graph in part a is like and unlike the graph of y=x2 at the beginning of this lesson

Answer by Alan3354(69443)

consider the formula A=s2 for the area A of a square with a side of length s.
a. graph all possible values of s and a on a coordinate plane.
b. explain how the graph in part a is like and unlike the graph of y=x2 at the beginning of this lesson
--------------
A = s^2 is the same as y = x^2. Only the labels are different.